Twenty years ago, when my parents were married, the average age of my parents was 23 years. After I was born, their average age was 34 years. What's the child's age?
Answer by richwmiller(17219)

(x-20+y-20)/2=23,
(x-z+y-z)/2=34
x+y-2z=68
x+y-40=46
-2z+40=22
20-z=11
z=9